x = 7 -2y
2x + y = 5
Solved by substitution or elimination

Answer by oscargut(2103)



x = 7 -2y
2x + y = 5
2(7-2y)+y = 5
14-4y+y = 5
-3y = -9
y = 3 , x=1
Answer: x= 1 , y = 3






You can ask me more at : mthman@gmail.com
Thanks